# Large Language Models Badly Generalize across Option Length, Problem Types, and Irrelevant Noun Replacements

### v2

@@ Abstract
In this paper, we propose a “Generalization Stress Test” to assess Large Language Models’ (LLMs) generalization ability under slight and controlled perturbations, including option length, problem types, and irrelevant noun replacements. We achieve novel and significant findings that, despite high benchmark scores, LLMs exhibit severe accuracy drops and unexpected biases (e.g., preference for longer distractors) when faced with these minor but content-preserving modifications. For example, Qwen 2.5 1.5B’s MMLU score rises from 60 to 89 and drops from 89 to 36 when option lengths are changed without altering the question. Even GPT4o experiences a 25-point accuracy loss when problem types are changed, with a 6-point drop across all three modification categories. These analyses suggest that LLMs rely heavily on superficial cues rather than forming robust, abstract representations that generalize across formats, lexical variations, and irrelevant content shifts.
@@ 1 Introduction
Recent work has focused on uncovering the actual limits of LLM generalization. One direction involves the development of dynamic evaluation methods that modify the evaluation process on the fly (Zhu et al., 2024; Yu et al., 2024). Another approach emphasizes creating more challenging or adversarial test sets that push models beyond their current capabilities, such as MMLU-Pro (Wang et al., 2024) and GSM-Plus (Li et al., 2024a). A third line of inquiry involves introducing subtle modifications to benchmark datasets to test LLM robustness, such as altering the order of multiple-choice options or changing the format of questions (Zheng et al., 2024; Li et al., 2024b; Gupta et al., 2024; Alzahrani et al., 2024). While these approaches have contributed to a better understanding of LLM performance, they either totally change the original problems, increase the complexity of the evaluation, or focus on relatively limited formatting changes like option ID adjustments.
@@ 3.2 Results of Altering Scoring Type
LLMs do not have invariant knowledge that can generalize across scoring types. As in Table 3, all models tend to score lower when the benchmarks are changed from the original format to boolean questions. Qwen2.5 1.5B and Llama3.1 8B score only half the points in the MMLU’s "both" setting. Smaller models generalize worse.
@@
Table 3: Performance on changing problem type from multi-choice question (MCQ) to bool questions (BQ). The values are percentages.
@@ 3.3 Results of Replacing Irrelevant Nouns
3.4 Reasoning Behind Accuracy Drops
The above analysis reveals that LLMs are severely biased to common but irrelevant patterns.
@@ 4 Conclusion
This paper finds that LLMs exhibit significant performance degradation when faced with slight changes in question format, option length, or irrelevant content shifts. These findings underscore that LLMs rely on superficial patterns rather than robust, generalizable reasoning. By introducing the "Generalization Stress Tests," we offer novel ideas towards evaluating LLMs’ true generalization capabilities.
@@
This work adheres to ACL’s ethical guidelines, we state that there are no ethical concerns to our knowledge.
@@ C.1 Evaluation Protocol
We perform evaluations on harness framework Gao et al. (2024). We chose harness because it is a flexible, configurable, reproducible framework. Unless otherwise specified, our evaluations are conducted in a 5-shot manner, with few-shot examples drawn from the benchmarks’ corresponding training sets.
@@ D.1 Making All Options longer
Benchmark Model Origin AL RL WL
MMLU Qwen2.5 1.5B 60.3 54.7 89.0 36.3
Qwen2.5 7B 73.7 69.2 90.1 55.6
Qwen2.5 72B 85.4 81.3 94.1 75.6
LLaMa3.1 8B 65.5 64.3 85.6 53.6
LLaMa3.1 70B 78.8 76.0 93.6 70.6
ARC-C Qwen2.5 1.5B 77.3 67.3 88.9 68.1
Qwen2.5 7B 90.0 85.3 94.3 84.0
Qwen2.5 72B 95.8 93.1 97.2 94.4
LLaMa3.1 8B 78.1 78.6 85.2 74.7
LLaMa3.1 70B 91.8 89.9 96.3 90.8

### v3

@@ Abstract
In this paper, we propose a “Generalization Stress Test” to assess Large Language Models’ (LLMs) generalization ability under slight and controlled perturbations, including option length, problem types, and irrelevant noun replacements. We achieve novel and significant findings that, despite high benchmark scores, LLMs exhibit severe accuracy drops and unexpected biases (e.g., preference for longer distractors) when faced with these minor but content-preserving modifications. For example, Qwen 2.5 1.5B’s MMLU score rises from 60 to 89 and drops from 89 to 36 when option lengths are changed without altering the question. Even GPT4o experiences a 25-point accuracy loss when problem types are changed, with a 6-point drop across all three modification categories. These analyses suggest that LLMs rely heavily on superficial cues rather than forming robust, abstract representations that generalize across formats, lexical variations, and irrelevant content shifts. Code can be found in: https://github.com/Qihoo360/LLMs-Generalization-Test.
@@ 1 Introduction
Recent work has focused on uncovering the actual limits of LLM generalization. One direction involves the development of dynamic evaluation methods that modify the evaluation process on the fly (Zhu et al., 2024; Yu et al., 2024), or extend the modality Wang et al. (2025). Another approach emphasizes creating more challenging or adversarial test sets that push models beyond their current capabilities, such as MMLU-Pro (Wang et al., 2024) and GSM-Plus (Li et al., 2024a). A third line of inquiry involves introducing subtle modifications to benchmark datasets to test LLM robustness, such as altering the order of multiple-choice options or changing the format of questions (Zheng et al., 2024; Li et al., 2024b; Gupta et al., 2024; Alzahrani et al., 2024; Hong et al., 2025). While these approaches have contributed to a better understanding of LLM performance, they either totally change the original problems, increase the complexity of the evaluation, or focus on relatively limited formatting changes like option ID adjustments.
@@ 3.2 Results of Altering Scoring Type
LLMs do not have invariant knowledge that can generalize across scoring types. As in Table 3, all models tend to score lower when the benchmarks are changed from the original format to boolean questions. Qwen2.5 1.5B and Llama3.1 8B score only half the points in the MMLU’s "both" setting. Smaller models generalize worse. [^3] The “MCQ” setting is equal to “Origin” setting in Table 1, the results are slightly different since we removed the instruction “Output the answer directly” to accommodate the BQ setting.
@@
Table 3: Performance on changing problem type from multi-choice question (MCQ) to bool questions (BQ). The values are percentages. “Both” means the percentages of examples whose MCQ and BQ are both true.
@@ 3.3 Results of Replacing Irrelevant Nouns
4 Discussion
4.1 Reasons Behind Accuracy Drops
The above ablation of results reveals that LLMs are severely biased to common but irrelevant patterns. Now, we delve a little deeper into root causes.
Could the imbalance in the test data be causing biased results? On the MMLU benchmark, a naïve policy that always selects the longest option reaches 28.3% accuracy, only +3.3% above the 25 % random baseline. This small gain shows that the length distribution of the options, by itself, is insufficient to yield the substantial performance gap reported in the paper.
Could the failures of large language models be attributed to certain mechanisms, such as the attention mechanism? Yes, we perform an analysis of attention patterns and find that lengthening options affects the attention mechanisms, causing LLMs to attend more to that option.
Condition A B C D
Origin 0.12 0.19 0.12 0.21
WL 0.10 0.16 0.26 0.17
Table 6: The attention scores of options. The scores are from layer 0 and are averaged across all 32 heads over 24 option-orders; the numbers are the summed attention weights on the choice tokens. A is the correct answer, while C is an intentionally lengthened distractor.
As in Table 6, the “WL” row shows that increasing the length of option C shifts more attention toward it, confirming a length-induced bias in the attention mechanism.
4.2 Generalization of Results
Could simple interventions (e.g., fine-tuning, CoT) address these vulnerabilities? We investigate both aspects by adding simple mitigations (fine-tuning) and Chain-of-Thought (CoT) prompting. On MMLU with Qwen2.5-7B, fine-tuning on augmented perturbations narrows robustness variance: WL improves by $7.8\%$ while RL decreases by $12.3\%$ versus the base model, partially mitigating length bias. CoT further shrinks the RL–WL gap from $34.5\%$ to $15.4\%$, but lowers overall accuracy ($10.7\%$), indicating that neither naive CoT nor simple fine-tuning fully resolves the vulnerability; stronger defenses (e.g., adversarial training or architectural changes) are likely required. CoT remains our default on GSM8K; here we additionally report MMLU results in Table 7.
Model Origin RL WL
Qwen2.5-7B 73.7 90.1 55.6
Qwen2.5-7B (fine-tuned) 69.1 77.8 63.4
Qwen2.5-7B (CoT) 63.0 71.7 56.3
Table 7: Accuracy (%) on MMLU for Qwen2.5-7B under simple mitigations: fine-tuning and Chain-of-Thought (CoT).
Could more proprietary LLMs, such as Gemini and Claude, be fragile to these changes?
Model Origin RL WL MCQ BQ
Qwen2.5 72B 85.4 94.1 75.6 84.0 69.1
GPT4o 85.2 89.7 83.3 84.7 59.5
Deepseek-v3 86.4 90.4 84.3 85.0 71.2
Claude-3.5-sonnet 86.1 92.2 85.6 87.9 30.6
Gemini-2.0-flash 85.8 90.3 85.0 86.1 41.3
Table 8: Accuracy (%) of SoTA short CoT models on MMLU.
As we can see from Table 8, SoTA short CoT models, including Gemini-2.0-flash [^4] Gemini 2: https://blog.google/technology/google-deepmind/google-gemini-ai-update-december-2024/, Claude-3.5-sonnet [^5] Claude 3.5 sonnet: https://www.anthropic.com/news/claude-3-5-sonnet, and Deepseek-v3 DeepSeek-AI (2025), are still sensitive to changes.
4.3 Compare to Work in the Pre-LLM Era
Indeed, early studies have shown the limited robustness of LLMs from 2018 to 2020 Naik et al. (2018); Zhang et al. (2019); Ribeiro et al. (2020). However, those analyses were limited to less than 1B-parameter models, whereas modern LLMs exhibit strong generalization abilities that may alter robustness patterns. Our work extends this line by examining SoTa models on comprehensive and premium-quality benchmarks and uncovers significant limitations of these capable LLMs.
4.4 Rank Changes of Models
As in D.3, the rank of models on the MMLU leaderboard changes across different settings.
@@ 5 Conclusion
This paper finds that LLMs exhibit significant performance degradation when faced with slight changes in question format, option length, or irrelevant content shifts. These findings underscore that LLMs rely on superficial patterns rather than robust, generalizable reasoning. By introducing the "Generalization Stress Tests," we offer novel understandings towards evaluating LLMs’ true generalization capabilities.
@@
This work adheres to ACL’s ethical guidelines, and we state that there are no ethical concerns to our knowledge.
Acknowledgments
We thank the anonymous reviewers, the AC, SAC, and PC for their contributions to this work, and we are grateful to Change Jia for involvement in the early exploration and to Hongye Zhang for the manual annotation.
@@ Appendix C Experiment Setup Details
Benchmark Model Origin AL RL WL
MMLU Qwen2.5 1.5B 60.3 54.7 89.0 36.3
Qwen2.5 7B 73.7 69.2 90.1 55.6
Qwen2.5 72B 85.4 81.3 94.1 75.6
LLaMa3.1 8B 65.5 64.3 85.6 53.6
LLaMa3.1 70B 78.8 76.0 93.6 70.6
ARC-C Qwen2.5 1.5B 77.3 67.3 88.9 68.1
Qwen2.5 7B 90.0 85.3 94.3 84.0
Qwen2.5 72B 95.8 93.1 97.2 94.4
LLaMa3.1 8B 78.1 78.6 85.2 74.7
LLaMa3.1 70B 91.8 89.9 96.3 90.8
@@ C.1 Evaluation Protocol
We perform evaluations on the Harness framework Gao et al. (2024). We chose Harness because it is a flexible, configurable, reproducible framework. Unless specified, we follow the default parameter of the harness. Unless otherwise specified, our evaluations are conducted in a 5-shot manner, with few-shot examples drawn from the benchmarks’ corresponding training sets.
@@ C.4 Budget
C.5 Random Seeds
All reported numbers are from a single wrong, since the model is deterministic with our default decoding temperature $T=0$, thus, changing seeds has no effect. We nevertheless tested robustness at $T=0.5$ and $T=1.0$ while varying Python/NumPy/Torch seeds; results for Qwen-2.5 7B on MMLU are shown below. The RL/WL gap persists, and our conclusions remain unchanged.
Temp. Seed Origin RL WL
0 0, 1234, 1234 73.7 90.1 55.7
1, 11, 111 73.7 90.1 55.7
2, 22, 222 73.7 90.1 55.7
0.5 0, 1234, 1234 71.6 88.2 54.9
1, 11, 111 71.5 88.5 54.9
2, 22, 222 71.5 88.3 55.3
1.0 0, 1234, 1234 67.9 83.4 52.2
1, 11, 111 67.3 83.8 51.9
2, 22, 222 67.2 83.8 53.0
Table 10: Model accuracy (%) for Qwen-2.5 7B on MMLU under different temperatures and seeds. Numbers are averaged over three runs when applicable; at $T=0$ the model is deterministic. Each seed setting contains three seeds regarding Python, NumPy, and Torch.
@@ D.2 Make All Wrong Options Longer
D.3 Rank Changes of Models
Table 12: Accuracy (%), rank, and rank changes of models on MMLU under different protocols.
Model Origin RL WL
Score Rank $\DeltaRank Score Rank Δ\DeltaRank$
Qwen2.5 1.5B 60.3 7 89.0 5  $\uparrow\,2$ 36.3 7
Qwen2.5 7B 73.7 5 90.1 3  $\uparrow\,2$ 55.6 5
Qwen2.5 72B 85.4 1 94.1 1 75.6 2  $\downarrow\,1$
LLaMa 3.1 8B 65.5 6 85.6 7  $\downarrow\,1$ 53.6 6
LLaMa 3.1 70B 78.8 3 93.6 2  $\uparrow\,1$ 70.6 3
GPT-4o mini 76.5 4 87.2 6  $\downarrow\,2$ 70.6 3  $\uparrow\,1$
GPT-4o 85.2 2 89.7 4  $\downarrow\,2$ 83.3 1  $\uparrow\,1$
Kendall $\tau$ 0.52 0.88
Table 13: Accuracy (%), rank, and rank changes of models on MMLU (MCQ vs. BQ).
Model MCQ BQ
[…]
Qwen2.5 1.5B 58.8 7 30.3 7
Qwen2.5 7B 72.4 5 54.7 4  $\uparrow\,1$
Qwen2.5 72B 84.0 2 69.1 1  $\uparrow\,1$
LLaMa 3.1 8B 64.6 6 40.6 6
LLaMa 3.1 70B 78.4 3 63.5 2  $\uparrow\,1$
GPT-4o mini 75.1 4 54.5 5  $\downarrow\,1$
GPT-4o 84.7 1 59.5 3  $\downarrow\,2$
Kendall $\tau$ 0.71
Note of Kendall $\tau$ . Kendall $\tau$ is a non-parametric rank-correlation coefficient that measures how similarly two lists are ordered by comparing the balance of concordant versus discordant item pairs, ranging from 1-1 (complete disagreement) to +1+1 (perfect agreement).
As we can see, the rank of the LLMs changes when we apply the perturbation. The Kendall $\tau$ between the WL and RL rankings is 0.39; since a Kendall $\tau$ below 0.4 is not considered strongly correlated, switching the evaluation protocol still has a pronounced influence on the rank shifts.